lopez spent 1/3 of his vacation money for travel and 2/5 of his vacation money for lodging. He spent $1100 for travel and lodging. What is the total amount of money he spent on his vacation?

Answer by rfer(16322)

1100/(2/5+1/3)=
1100/(6/15+5/15)=
1100/11/15=
1100*15/11=$1500